Determine the area of the surface generated by revolving a given curve about a given axis.

a.) y=x³ on (0,1), about the x-axis.

b. ) y=x² on (0,√2), about the y-axis.


1
Expert's answer
2021-05-31T03:43:02-0400

a)

"y=x^3"

"\frac{dy}{dx}=3x^2"

"Surface\ area\ of\ revolution"

"2\pi\int_{0}^{1}{x^3\sqrt{1+\left(3x^2\right)^2}\ dx}"

"let\ u=1+9x^4"

"du=36x^3\ dx"

"2\pi\int{\sqrt u\frac{du}{36}}"

"\frac{\pi}{18}\left(\frac{2}{3}\right)\left(u\right)^\frac{3}{2}"

"\frac{\pi}{27}\left(1+9x^4\right)^\frac{3}{2}\ \ \ \ \ \left(0,\ 1\right)"

"\frac{\pi}{27}\left(\left(10\right)^\frac{3}{2}-1\right)"

"1.13412\pi"


b)

"y=x^2"

"x=\ \sqrt y"

"\frac{dx}{dy}=\frac{1}{2\sqrt y}"

"Surface\ are\ of\ revolution"

"2\pi\int_{0}^{\sqrt2}{\sqrt y\sqrt{1+\left(\frac{1}{2\sqrt y}\right)^2}\ dy}"

"2\pi\int_{0}^{\sqrt2}{\sqrt{y+\frac{1}{4}}\ dy}"

"2\pi\ \left(\frac{2}{3}\right)\left(y+\frac{1}{4}\right)^\frac{3}{2}\ (0,\ \sqrt2)"

"\frac{4\pi\ }{3}\left({\ \left(\sqrt2+\frac{1}{4}\right)}^\frac{3}{2}-\left(\frac{1}{4}\right)^\frac{3}{2}\right)"

"2.6958\pi\"